# Thorough Analysis of $\pi^{-}p\rightarrow K^{*}\Sigma$ Reaction: Probing Nucleon Excited State and Constituent Scaling Laws

### v1

@@ V Dalitz process and experimental feasibility
In the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the $K^{*}$ meson produced cannot be measured directly. Instead, it must be determined by reconstructing the final particles resulting from its decay. Given that the branching ratio for $K^{*}$ decay to $K^{+}\pi^{-}$ is as high as $99.754\%$ ParticleDataGroup:2022pth , analyzing the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ is crucial for providing useful information for experimental measurements. The basic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ reaction are illustrated in Fig. 16. Our calculations show that the total cross section for $N^{*}(2080)(\frac{1}{2}^{-})$ is slightly larger than that for the $N^{*}(2080)(\frac{3}{2}^{-})$. In this study, we focus on the Dalitz process specifically for $N^{*}$ with $J^{P}=\frac{3}{2}^{-}$. Generally, the invariant mass spectrum of the Dalitz process is defined based on the two-body process Kim:2017nxg ; Wang:2024qnk
@@
The calculated total cross sections for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and the experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are presented in Table 6. To assess the feasibility of experimentally detecting the $K^{*}$ via $\pi^{-}p$ scattering, we calculate the ratio $\sigma(\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0})/\sigma(\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0})$. The minimum ratio reaches 29.67% at center-of-mass energies between 2.56 and 2.88 GeV, indicating that the total cross section and event number for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process reconstructed via $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are sufficient for experimental measurements. Additionally, the ratio can reach up to 76.61% at a center-of-mass energy of 2.56 GeV, with the $s$-channel $N^{*}(2080)$ exchange contributing 73.19%, approximately 30.28 $\mu\text{b}$. These results suggest that a significant number of the $N^{*}(2080)$ can be produced in the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process, providing strong support for future experimental efforts.
@@
As a leading hadron experimental facility, J-PARC currently provides the highest-intensity meson beams (both $K$ and $\pi$ mesons) in the several GeV energy range, making it an outstanding platform for particle and nuclear physics research Aoki:2021cqa . Theoretical studies and experimental evidence suggest that certain light-flavored baryons may contain a significant strange quark pair ($s\bar{s}$) component, potentially forming a hidden-strange pentaquark state, such as the $N^{*}(2080)$. However, the existence of a hidden-strange pentaquark state has not yet been confirmed experimentally, underscoring the need for collaborative efforts between experimental and theoretical research. If J-PARC can measure the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the findings could aid in the discovery of the hidden-strange pentaquark state $N^{*}(2080)$ and provide valuable insights into quark interactions within hadrons, particularly the internal dynamics of diquarks.
@@
Additionally, HIAF offers high-intensity proton and ion beam currents across a wide energy range. Upon completion, HIAF will deliver the highest peak-intensity medium and low-energy heavy-ion beam currents compared to similar facilities worldwide HIAF . If a meson beam experimental device is developed at HIAF in the future, it could enable high-precision measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, further supporting the search for the $N^{*}(2080)$ wl:2024ext .
VI Summary and prospect
@@
In recent years, the $N^{*}(2080)$ has been considered the hidden-strange partner of the hidden-charm molecular pentaquark $P_{c}(4457)$, as its mass lies just below the $K^{*}\Sigma$ threshold He:2017aps ; Ben:2023uev ; Xie:2010yk ; Lin:2018kcc . In this study, we investigate the production of the nucleon excited state $N^{*}(2080)$ through the $\pi^{-}p$ scattering process using the effective Lagrangian approach. Specifically, we consider $K$ and $K^{*}$ meson exchanges in the $t$-channel, $\Sigma$ exchange in the $u$-channel, and the contribution of the $N^{*}(2080)$ in the $s$-channel. By fitting the total and differential cross sections of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we demonstrate that the $N^{*}(2080)$ exchange provides the dominant contribution to the total cross section near the threshold energy region, offering potential insights into detecting the $N^{*}(2080)$ through the $\pi^{-}p$ scattering process. However, the lack of experimental data near the threshold imposes limitations on determining the properties of the $N^{*}(2080)$ via this process.
@@
Given the significant branching ratio of $K^{*}$ decay into $K^{+}\pi^{-}$, we calculate the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and compare it with experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$. The results suggest that it is feasible to observe the $N^{*}(2080)$ through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process in experiments. Notably, our calculations consider two scenarios for the $J^{P}$ quantum number of the $N^{*}(2080)$, either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$. A significant source of uncertainty in our model arises from the absence of experimental measurements of the partial widths of the $N^{*}(2080)$ decays to $\pi N$, leading us to treat the coupling constants $g_{\pi NN^{*}}^{3/2^{-}}$ and $g_{\pi NN^{*}}^{1/2^{-}}$ as free parameters determined by fitting. In conclusion, further experimental measurements and theoretical investigations into the internal properties of $N^{*}$ states are highly desirable.
Additionally, the application of the constituent counting rule in hard exclusive reactions is discussed, including a demonstration of the rule by deriving expressions for the reaction cross-section and matrix element, as well as explaining the contribution of typical hard gluon exchange processes to exclusive reactions. By calculating the mass dimension, we derive the constituent counting expression for the reaction cross-section. In the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, the predictions of this rule align with experimental data, further validating its accuracy.
@@
Currently, high-precision measurements of the meson-nucleus scattering process can be conducted by experiments at J-PARC Aoki:2021cqa , AMBER Adams:2018pwt , and future facilities such as HIKE HIKE:2023ext and HIAF HIAF . We recommend that these experiments perform precise measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, particularly the differential cross section at the forward angle ($\theta=0^{\circ}$) and the $t$-distribution cross section at large momentum transfer ($\theta=90^{\circ}$). This is crucial for elucidating the mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process and determining the contribution of the $N^{*}$ state. To date, pentaquark-like states containing strange quarks have not been experimentally confirmed, necessitating further clarification through collaborative experimental and theoretical efforts. Analyzing the existence and contribution of pentaquark-like states with strange quarks via the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction is an effective approach wl:2024ext . Moving forward, we plan to systematically investigate the contribution of excited baryon states in various $\pi^{-}p$ or $K^{-}p$ scattering processes, providing essential theoretical support for future experiments.

### v2

@@ IV Dalitz process and experimental feasibility
In the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the $K^{*}$ meson produced cannot be measured directly. Instead, it must be determined by reconstructing the final particles resulting from its decay. Given that the branching ratio for $K^{*}$ decay to $K^{+}\pi^{-}$ is as high as $99.754\%$ ParticleDataGroup:2022pth , analyzing the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ is crucial for providing useful information for experimental measurements. The basic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ reaction are illustrated in Fig. 16. Our calculations show that the total cross section for $P_{s\bar{s}}[1/2^{-}]$ is slightly larger than that for the $P_{s\bar{s}}[3/2^{-}]$. In this study, we focus on the Dalitz process specifically for $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$. Generally, the invariant mass spectrum of the Dalitz process is defined based on the two-body process Kim:2017nxg ; Wang:2024qnk
@@
The calculated total cross sections for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and the experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are presented in Table 6. To assess the feasibility of experimentally detecting the $K^{*}$ via $\pi^{-}p$ scattering, we calculate the ratio $\sigma(\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0})/\sigma(\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0})$. The minimum ratio reaches 29.67% at center-of-mass energies between 2.56 and 2.88 GeV, indicating that the total cross section and event number for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process reconstructed via $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are sufficient for experimental measurements. Additionally, the ratio can reach up to 76.61% at a center-of-mass energy of 2.56 GeV, with the $s$-channel $P_{s\bar{s}}$ exchange contributing 73.19%, approximately 30.28 $\mu\text{b}$. These results suggest that a significant number of the $P_{s\bar{s}}$ can be produced in the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process, providing strong support for future experimental efforts.
@@
As a leading hadron experimental facility, J-PARC currently provides the highest-intensity meson beams (both $K$ and $\pi$ mesons) in the several GeV energy range, making it an outstanding platform for particle and nuclear physics research Aoki:2021cqa . Theoretical studies and experimental evidence suggest that certain light-flavored baryons may contain a significant strange quark pair ($s\bar{s}$) component, potentially forming a hidden-strange pentaquark state, such as the $P_{s\bar{s}}$. However, the existence of a hidden-strange pentaquark state has not yet been confirmed experimentally, underscoring the need for collaborative efforts between experimental and theoretical research. If J-PARC can measure the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the findings could aid in the discovery of the hidden-strange pentaquark state $P_{s\bar{s}}$ and provide valuable insights into quark interactions within hadrons, particularly the internal dynamics of diquarks.
@@
Additionally, HIAF offers high-intensity proton and ion beam currents across a wide energy range. Upon completion, HIAF will deliver the highest peak-intensity medium and low-energy heavy-ion beam currents compared to similar facilities worldwide HIAF . If a meson beam experimental device is developed at HIAF in the future, it could enable high-precision measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, further supporting the search for the $P_{s\bar{s}}$ wl:2024ext .
V Summary and prospect
@@
In recent years, the $P_{s\bar{s}}$ has been considered the hidden-strange partner of the hidden-charm molecular pentaquark $P_{c}(4457)$, as its mass lies just below the $K^{*}\Sigma$ threshold He:2017aps ; Ben:2023uev ; Xie:2010yk ; Lin:2018kcc . In this study, we investigate the production of the hidden-strange molecular-type pentaquarks $P_{s\bar{s}}$ through the $\pi^{-}p$ scattering process using the effective Lagrangian approach. Specifically, we consider $K$ and $K^{*}$ meson exchanges in the $t$-channel, $\Sigma$ exchange in the $u$-channel, and the contribution of the $P_{s\bar{s}}$ in the $s$-channel. By fitting the total and differential cross sections of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we demonstrate that the $P_{s\bar{s}}$ exchange provides the dominant contribution to the total cross section near the threshold energy region, offering potential insights into detecting the $P_{s\bar{s}}$ through the $\pi^{-}p$ scattering process. However, the lack of experimental data near the threshold imposes limitations on determining the properties of the $P_{s\bar{s}}$ via this process.
@@
Given the significant branching ratio of $K^{*}$ decay into $K^{+}\pi^{-}$, we calculate the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and compare it with experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$. The results suggest that it is feasible to observe the $P_{s\bar{s}}$ through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process in experiments. Notably, our calculations consider two scenarios for the $J^{P}$ quantum number of the $P_{s\bar{s}}$, either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$. A major source of uncertainty in our model stems from the lack of experimental data on the partial widths of $P_{s\bar{s}}$ decays to $\pi N$. This necessitates treating the coupling constants $g_{\pi NP^{*}}^{3/2^{-}}$ and $g_{\pi NP^{*}}^{1/2^{-}}$ as free parameters, determined through fitting. In these fitting calculations, both the coupling constants and the cutoff parameters in the form factors are set as free parameters, which is an inherent aspect of this theoretical approach. Consequently, while we can calculate and analyze the production mechanism and cross section of the $P_{s\bar{s}}$ state via the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we cannot further deduce the internal structural properties of the $P_{s\bar{s}}$ state. Additionally, the application of the constituent counting rule in hard exclusive reactions is discussed. For the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, our theoretical cross section predictions align well with the constituent counting rules, providing a strong validation of the overall cross-section calculation.
@@
Currently, high-precision measurements of the meson-nucleus scattering process can be conducted by experiments at J-PARC Aoki:2021cqa , AMBER Adams:2018pwt , and future facilities such as HIKE HIKE:2023ext and HIAF HIAF . We recommend that these experiments perform precise measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, particularly the differential cross section at the forward angle ($\theta=0^{\circ}$) and the $t$-distribution cross section at large momentum transfer ($\theta=90^{\circ}$). This is crucial for elucidating the mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process and determining the contribution of the $P_{s\bar{s}}$ state. To date, pentaquark-like states containing strange quarks have not been experimentally confirmed, necessitating further clarification through collaborative experimental and theoretical efforts. Analyzing the existence and contribution of pentaquark-like states with strange quarks via the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction is an effective approach wl:2024ext . Moving forward, we plan to systematically investigate the contribution of excited baryon states in various $\pi^{-}p$ or $K^{-}p$ scattering processes, providing essential theoretical support for future experiments.